ABCcompany is considering a capital investment for which initial outlay is 500,000 and has estimated life of five years with a zero salvage value, ignoring depreciation, the machine produced following cash flows:- in year 1, $70,000, in year 2, 60,000, in year 3, 50,000, in year 4, 60,000, in year 5, 70,000 compute NPV and pay back period, discount rate is 8%, interprete your finding.
1
Expert's answer
2015-02-26T09:28:06-0500
I = 500,000, cash flows: in year 1, $70,000, in year 2, 60,000, in year 3, 50,000, in year 4, 60,000, in year 5, 70,000, r = 0.08
NPV = -I + CF1/(1 + r)^1 +...+ CF1/(1 + r)^n = -500,000 + 70,000/1.08 + 60,000/1.08^2 + 50,000/1.08^3 + 60,000/1.08^4 + 70,000/1.08^5 = -$252,310.63 and pay back period is more than 5 years.
So, this investment will not be paid back and is not efficient.